OpenStudy (anonymous):

IDENTIFY THE SLOPE INTERCEPT FOMS OF 4y=x+8

OpenStudy (anonymous):

y=1/4x + 2

OpenStudy (anonymous):

is this in slope intercept form

OpenStudy (anonymous):

Yes slope intercept is y=mx+b

OpenStudy (anonymous):

no the question says write 4y=x+8 in slope intercept form

OpenStudy (anonymous):

Yes and y=1/4x + 2 is 4y=x+8 in slope intercept form.

OpenStudy (anonymous):

3y+2x=6